# LazyDrag: Enabling Stable Drag-Based Editing on Multi-Modal Diffusion Transformers via Explicit Correspondence

### v2

@@ 3.3.1. Attention Input Control via Token Replacement and Concatenation
For the background region $\mathcal{R}^{\text{bg}}$, the purpose of absolute untouched is achieved by hard-replacing the attention tokens with their cached originals at every step and every single-stream layer, similar to ColorCtrl (Yin et al., 2025):
(3) $(\mathbf{Q}_{\bm{x}},\mathbf{K}_{\bm{x}},\mathbf{V}_{\bm{x}})\leftarrow(\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{Q}}_{\bm{x}}),\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{K}}_{\bm{x}}),\overline{\mathbf{V}}_{\bm{x}}),\quad\forall\bm{x}\in\mathcal{R}^{\text{bg}}.$
@@
(4) $\tilde{\mathcal{M}}(\bm{x})=\begin{cases}\mathcal{M}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\text{dst}},\\ \bm{x},&\text{if }\bm{x}\in\mathcal{R}^{\text{trans}}.\end{cases}$
@@
(5) $\mathbf{K}^{\prime}_{\bm{x}}$ $=\operatorname{concat}\big(\mathbf{K}_{\bm{x}},\,\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{K}}_{\tilde{\mathcal{M}}(\bm{x})})\big),$
(6) $\mathbf{V}^{\prime}_{\bm{x}}$ $=\operatorname{concat}\big(\mathbf{V}_{\bm{x}},\,\overline{\mathbf{V}}_{\tilde{\mathcal{M}}(\bm{x})}\big).$
@@
3.3.2. Attention Output Refinement via Gated Merging (Attn Refine)
The second part refines the attention output so that it cooperates with the above token concatenation (following Wang et al. (2025a)), improving visual quality and emphasizing the importance of handle points over others. Let $\mathbf{y}_{\bm{x}}$ be the attention output at $\bm{x}$ and $\overline{\mathbf{y}}_{\bm{x}}$ be the cached output. For $\bm{x}\in\mathcal{R}^{\mathrm{dst}}$,
(7) $\mathbf{y}_{\bm{x}}\leftarrow\bigl(1-\gamma_{\bm{x},t}\bigr)\,\mathbf{y}_{\bm{x}}+\gamma_{\bm{x},t}\,\overline{\mathbf{y}}_{\mathcal{M}(\bm{x})},$
@@ 3.3.2. Attention Output Refinement via Gated Merging (Attn Refine)
(8) $\gamma_{\bm{x},t}=h_{t}\cdot\mathcal{A}(\bm{x}),$
where $t$ indexes the timestep and $h_{t}\in[0,1]$ is a factor that decays over time. This correspondence-driven gated merge eliminating the extra denoising steps required by CharaConsist, and addressing the instability of attention-similarity matching and scaling under full-strength inversion. By making the weight strongest at the handle points (where $\mathcal{A}(\bm{x})$ is maximal), it ensures precise control where it matters most, removing the need for multi-step latent optimization in previous methods (Zhang et al., 2025; Shi et al., 2024), while allowing for natural relaxation in surrounding regions.
4. Experiments
4.1. Setup
@@
We compare against eight prior arts: DragDiffusion (Shi et al., 2024), DragNoise (Liu et al., 2024a), FreeDrag (Ling et al., 2024), DiffEditor (Mou et al., 2024a), GoodDrag (Zhang et al., 2025), DragText (Choi et al., 2025) [^1] Since DragText is a plug-and-play method, we evaluate it in conjunction with best-performing GoodDrag., FastDrag (Zhao et al., 2024), and Inpaint4Drag (Lu and Han, 2025). Others are excluded due to incompatible settings or unusable official implementations.
@@
Unless otherwise noted, all baselines are run with their official implementations and default hyperparameters. For Inpaint4Drag (Lu and Han, 2025), we adopt the refined masks and point pairs provided by the authors at inference, and replace distilled models with original models. Our method is built on FLUX.1 Krea-dev (Labs, 2025), adopting the inversion method of UniEdit-Flow (Jiao et al., 2025) while replacing the editing strategy with our approach. For a fair comparison, the number of denoising steps is fixed to 50 for all methods. Following CharaConsist, we activate ID Pres. and Attn Refine (Sec. 3.3) for the first 40 denoising steps, referring to the last activate timestep as the activation timestep. Additional details are in Appendix A.1.
@@
We evaluate on DragBench (Shi et al., 2024), which contains 205 images with 349 handle and target point pairs. Our primary accuracy metric is MD (mean distance) (Pan et al., 2023). Although IF (image fidelity) (Kawar et al., 2023), typically computed with LPIPS (Zhang et al., 2018), is widely used, we do not report IF. Previous work (Choi et al., 2025; Lu et al., 2024) shows that successful drag edits necessarily change the image, often increasing LPIPS, whereas an unchanged image trivially attains the best score. Hence, IF can be misleading for drag editing. To obtain a complementary, perceptually grounded view, we adopt the VIEScore (Ku et al., 2024) metrics from GEdit-Bench (Liu et al., 2025a): SC (Semantic Consistency): whether the intended edit has been achieved. PQ (Perceptual Quality): the naturalness of the result and absence of artifacts. O (Overall): the overall performance defined in Liu et al. (2025a). In our setting, the “intended edit” is specified by the dragging instruction rather than a natural-language instruction, but the scoring criteria remain unchanged. Each score ranges from 0 to 10 (higher is better) and is produced by the state-of-the-art MLLM evaluator, GPT-4o [^2] API access as of August 2025 (Hurst et al., 2024). To mitigate stochasticity in evaluation, we run every evaluation metrics three times and report both the mean and standard deviation. We additionally report a binary TTO-Req (Test-Time Optimization Required) flag indicating whether a method requires per-edit test-time optimization (e.g., LoRA fine-tuning or multi-step latent optimization) during inference. More evaluation details are in Appendix A.3.
@@
Table 1. Quantitative results compared with baselines on Drag-Bench.
Figure 4. Qualitative results compared with baselines on Drag-Bench. Best viewed with zoom-in.
4.2. Quantitative Evaluation
@@
4.3. Qualitative Evaluation
Fig. 4 qualitatively demonstrates the superiority of our method over existing baselines. In the first example, only our method correctly lift the arm with background maintained, while others introduce artifacts, such as distorted hands (e.g., DragText (Choi et al., 2025)) or unintended background changes (e.g., DragNoise (Liu et al., 2024a)). In the second example, most baselines fail to preserve the front structure of the vehicle, whereas our approach maintains it faithfully while applying the desired transformation. Inpaint4Drag (Lu and Han, 2025) generates artifacts in the background. The third case shows that only our method successfully modifies the sofa geometry while preserving the integrity of pillows. In the fourth example, our approach correctly interprets hand proximity as intent to insert it into the pocket, while other baselines introducing artifacts. Finally, in the fifth example, only our approach and DragText successfully rotates the head of the tiger to the right without compromising overall image quality. These results are consistent with our quantitative evaluations and highlight the robustness and generality of our method, even without per image tuning or per instruction multi-step latent optimization. More results are shown in Appendix B.1.
4.4. User Study
@@
DragNoise (Liu et al., 2024a) 5.00 $\pm$ 7
DragDiffusion (Shi et al., 2024) 8.75 $\pm$ 8
FreeDrag (Ling et al., 2024) 8.75 $\pm$ 8
DiffEditor (Mou et al., 2024a) 2.50 $\pm$ 4
GoodDrag (Zhang et al., 2025) 3.75 $\pm$ 4
DragText (Choi et al., 2025) 3.13 $\pm$ 4
FastDrag (Zhao et al., 2024) 3.13 $\pm$ 6
@@
Ours 61.88 $\pm$ 19
Table 2. User study on Drag-Bench.
Figure 5. Comparison between drag and move mode on Drag-Bench.
A total of 20 expert participants evaluated comparisons between methods on 32 cases randomly sampled from DragBench. For each comparison, method order positions were randomized and method identities were anonymized. Participants selected the preferred result according to predefined criteria (edit success, naturalness, and background preservation). Overall, LazyDrag was preferred in 61.88% of comparisons, outperforming all baselines (Tab. 2). More details are in Appendix A.4.
Figure 6. Qualitative cumulative ablation on Drag-Bench. Rows remove one component relative to the row above. When WTA and Latent Init are removed we use latent init in FastDrag. When ID Pres. and Attn Refine are removed we switch to CharaConsist attention-similarity control.
4.5. Comparison Between Drag and Move Modes
@@ 4.6. Ablation Study
We conduct an ablation study in which components are progressively removed from the full method. To keep functionality comparable when a component is absent, we adopt controlled replacements: (i) Without WTA and Latent Init (Sec. 3.2) we revert to latent warpage optimization of FastDrag (Zhao et al., 2024) as the latent initialization. (ii) Without ID Pres. and Attn Refine (Sec. 3.3) we switch to the attention-similarity matching and scaling introduced in CharaConsist (Wang et al., 2025a). Fig. 6 and Tab. 3 report benchmark results on Drag-Bench. Removing WTA and Latent Init increases MD and slightly reduces PQ and O, indicating that our initialization with the winner-takes-all fusion strategy and random initialization for inpainting regions suppresses repetitive artifacts and improves inpainting quality as proven in the figure. Further disabling background preservation causes additional drops in SC and O due to color shifting and artifacts in the background. Finally, replacing our correspondence-driven preservation with attention-similarity control leads to a sharp degradation, highlighting the sensitivity of full-strength inversion to mismatched attention alignment. The full method achieves the best performance.
@@
We conduct an ablation study on the effect of activation timesteps by varying the activation timestep to 20, 40, and 50, as shown in Fig. 7 and Tab. 4. From the results, we observe that increasing the number of the activation timestep leads to more accurate destination points for dragging, though it may introduce more warping artifacts. Conversely, reducing the activation timestep results in more natural outputs, but may cause slight variations in identity or motion. More results are in Appendix B.4. For benchmark evaluations, we use 40 as a balanced value.
Figure 7. Qualitative ablation of activation timesteps on Drag-Bench. From left to right, the activation timestep is increased.
@@
Table 3. Quantitative cumulative ablation on Drag-Bench under the same setting as Fig. 6
@@
Table 4. Quantitative ablation of activation timesteps on Drag-Bench.
5. Conclusion
@@ A.1. Inference Settings
EasyDrag (Hou et al., 2024) and CLIPDrag (Jiang et al., 2025) are excluded from comparison because their released implementations either fail to execute reliably or do not reproduce the results reported in the papers.
@@
For our inversion process, we adopt the official inversion method of UniEdit-Flow (Jiao et al., 2025) but replace the editing component with our proposed strategy. We apply our correspondence-driven preservation (Sec. 3.3) only to the single-stream attention layers in FLUX.1 Krea-dev (Labs, 2025). Since additional manipulation in dual-stream attention layers does not lead to noticeable improvements (Deng et al., 2025; Yin et al., 2025; Wang et al., 2025b), we adopt a more efficient and concise design by limiting modifications to single-stream layers only.
A.2. Implementation Details of Displacement Field Calculation
@@
(10) $\lambda_{j}^{i}=\frac{\lVert\bm{v}_{j}^{i}\rVert_{2}}{\lVert\bm{d}_{i}\rVert_{2}}=\frac{\lVert\bm{p}_{j}-\bm{p}_{j}^{i}\rVert_{2}}{\lVert\bm{s}_{i}-\bm{e}_{i}\rVert_{2}}=\frac{\lVert\bm{p}_{j}-\bm{q}_{j}^{i}\rVert_{2}}{\lVert\bm{s}_{i}-\bm{q}_{j}^{i}\rVert_{2}}.$
@@
(11) $\alpha_{j}^{i}=\begin{cases}\lVert\bm{p}_{j}-\bm{s}_{i}\rVert_{2}^{-1},&\bm{p}_{j}\neq\bm{s}_{i},\\ \infty,&\text{otherwise}.\end{cases}$
@@
(12) $\bm{v}_{j}=\bm{v}_{j}^{i^{\star}}=\lambda_{j}^{i^{\star}}\,\bm{d}_{i^{\star}}.$
@@
(13) $\bm{v}_{j}=\lambda_{j}^{i^{\star}}\,\bm{d}_{i^{\star}}+(\bm{r}-\mathbf{1})\otimes(\bm{p}_{j}-s_{i^{\star}}),$
@@
A.3. Evaluation Details
For the VIEScore evaluation, we follow GEdit-Bench (Liu et al., 2025a), using the same prompts for PQ and O. For SC, we adopt the instruction shown in Fig. 14, together with the source image, drag-instruction image, and the edited image. Score collection and calculation are carried out using the official GEdit-Bench codebase.
A.4. User Study Details
To evaluate the effectiveness of our method, we randomly selected 32 results for nine comparison methods on Drag-Bench (Shi et al., 2024) and shuffled their indices to ensure a fair comparison. We invited 20 participants, each with relevant skills, to perform the tasks following the instructions provided through the user interface, as shown in Fig. 13.

### v3

@@ 3.3.1 Attention Input Control via Token Replacement and Concatenation
For the background region $\mathcal{R}^{\text{bg}}$, the purpose of absolute untouched is achieved by hard-replacing the attention tokens with their cached originals at every step and every single-stream layer, similar to ColorCtrl (Yin et al., 2025b):
$(\mathbf{Q}_{\bm{x}},\mathbf{K}_{\bm{x}},\mathbf{V}_{\bm{x}})\leftarrow(\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{Q}}_{\bm{x}}),\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{K}}_{\bm{x}}),\overline{\mathbf{V}}_{\bm{x}}),\quad\forall\bm{x}\in\mathcal{R}^{\text{bg}}.$ (3)
@@
$\tilde{\mathcal{M}}(\bm{x})=\begin{cases}\mathcal{M}(\bm{x}),&\text{if }\bm{x}\in\mathcal{R}^{\text{dst}},\\ \bm{x},&\text{if }\bm{x}\in\mathcal{R}^{\text{trans}}.\end{cases}$ (4)
@@
$\mathbf{K}^{\prime}_{\bm{x}}$ $=\operatorname{concat}\big(\mathbf{K}_{\bm{x}},\,\mathrm{RoPE}_{\bm{x}}(\overline{\mathbf{K}}_{\tilde{\mathcal{M}}(\bm{x})})\big),$ (5)
$\mathbf{V}^{\prime}_{\bm{x}}$ $=\operatorname{concat}\big(\mathbf{V}_{\bm{x}},\,\overline{\mathbf{V}}_{\tilde{\mathcal{M}}(\bm{x})}\big).$ (6)
@@
3.3.2 Attention Output Refinement via Gated Merging (Attn Refine)
The second part refines the attention output so that it cooperates with the above token concatenation (following Wang et al. (2025b)), improving visual quality and emphasizing the importance of handle points over others. Let $\mathbf{y}_{\bm{x}}$ be the attention output at $\bm{x}$ and $\overline{\mathbf{y}}_{\bm{x}}$ be the cached output. For $\bm{x}\in\mathcal{R}^{\mathrm{dst}}$,
$\mathbf{y}_{\bm{x}}\leftarrow\bigl(1-\gamma_{\bm{x},t}\bigr)\,\mathbf{y}_{\bm{x}}+\gamma_{\bm{x},t}\,\overline{\mathbf{y}}_{\mathcal{M}(\bm{x})},$ (7)
@@ 3.3.2 Attention Output Refinement via Gated Merging (Attn Refine)
$\gamma_{\bm{x},t}=h_{t}\cdot\mathcal{A}(\bm{x}),$ (8)
where $t$ indexes the timestep and $h_{t}\in[0,1]$ is a factor that decays over time. This correspondence-driven gated merge eliminating the extra denoising steps required by CharaConsist, and addressing the instability of attention-similarity matching and scaling under full-strength inversion. By making the weight strongest at the handle points (where $\mathcal{A}(\bm{x})$ is maximal), it ensures precise control where it matters most, removing the need for multi-step latent optimization in previous methods (Zhang et al., 2025; Shi et al., 2024b), while allowing for natural relaxation in surrounding regions.
4 Experiments
4.1 Setup
@@
We compare against eight baselines: DragDiffusion (Shi et al., 2024b), DragNoise (Liu et al., 2024a), FreeDrag (Ling et al., 2024), DiffEditor (Mou et al., 2024a), GoodDrag (Zhang et al., 2025), DragText (Choi et al., 2025) [^1] Since DragText is a plug-and-play method, we evaluate it in conjunction with best-performing GoodDrag., FastDrag (Zhao et al., 2024), and Inpaint4Drag (Lu and Han, 2025). Notably, all baselines are U-Net–based, whereas ours is the first MM-DiT-based method.
@@
Unless otherwise noted, all baselines are run with their official implementations and default hyperparameters. For Inpaint4Drag (Lu and Han, 2025), we adopt the refined masks and point pairs provided by the authors at inference, and replace distilled models with original models. Our method is built on FLUX.1 Krea-dev (Labs, 2025), adopting the inversion method of UniEdit-Flow (Jiao et al., 2025) while replacing the editing strategy with our approach. Following CharaConsist (Wang et al., 2025b), we activate ID Pres. and Attn Refine (Sec. 3.3) for the first 40 denoising steps, referring to the last activate timestep as the activation timestep. For a fair comparison, the number of denoising steps is fixed to 50 for all methods. More details are in Appendix A.1.
@@
We evaluate on DragBench (Shi et al., 2024b), which contains 205 images with 349 handle and target point pairs. Our primary accuracy metric is MD (mean distance) (Pan et al., 2023). Although IF (image fidelity) (Kawar et al., 2023), typically computed with LPIPS (Zhang et al., 2018), is widely used, we do not report IF. Previous work (Choi et al., 2025; Lu et al., 2024) shows that successful drag edits necessarily change the image, often increasing LPIPS, whereas an unchanged image trivially attains the best score. Hence, IF can be misleading for drag editing. To obtain a complementary, perceptually grounded view, we adopt the VIEScore (Ku et al., 2024) metrics from GEdit-Bench (Liu et al., 2025b): SC (Semantic Consistency): whether the intended edit has been achieved. PQ (Perceptual Quality): the naturalness of the result and absence of artifacts. O (Overall): the overall performance defined in Liu et al. (2025b). In our setting, the “intended edit” is specified by the dragging instruction rather than a natural-language instruction, but the scoring criteria remain unchanged. Each score ranges from 0 to 10 (higher is better) and is produced by the state-of-the-art MLLM evaluator, GPT-4o [^2] API access as of August 2025 (Hurst et al., 2024). To mitigate stochasticity in evaluation, we run every evaluation metrics three times and report both the mean and standard deviation. We additionally report a binary TTO-Req (Test-Time Optimization Required) flag indicating whether a method requires per-edit test-time optimization (e.g., LoRA fine-tuning or multi-step latent optimization) during inference. More evaluation details are in Appendix A.3.
Table 1: Quantitative results compared with baselines on Drag-Bench.
@@
Figure 4: Qualitative results compared with baselines on Drag-Bench. Best viewed with zoom-in.
4.2 Quantitative Evaluation
@@
4.3 Qualitative Evaluation
Fig. 4 qualitatively demonstrates the superiority of our method over existing baselines. In the first example, only our method correctly lift the arm with background maintained, while others introduce artifacts, such as distorted hands (e.g., DragText (Choi et al., 2025)) or unintended background changes (e.g., DragNoise (Liu et al., 2024a)). In the second example, most baselines fail to preserve the front structure of the vehicle, whereas our approach maintains it faithfully while applying the desired transformation. Specifically, Inpaint4Drag (Lu and Han, 2025) generates artifacts in the background. The third case shows that only our method successfully modifies the sofa geometry while preserving the integrity of pillows. In the fourth example, our approach correctly interprets hand proximity as intent to insert it into the pocket, while other baselines introducing artifacts. Finally, in the fifth example, only our approach and DragText successfully rotates the head of the tiger to the right without compromising overall image quality. These results are consistent with our quantitative evaluations and highlight the robustness and generality of our method, even without per image tuning or per instruction multi-step latent optimization. More results are shown in Appendix B.
4.4 User Study
Table 2: User study on Drag-Bench.
@@ 4.4 User Study
DragNoise (Liu et al., 2024a) 6.64 $\pm$ 8
DragDiffusion (Shi et al., 2024b) 6.25 $\pm$ 8
FreeDrag (Ling et al., 2024) 6.64 $\pm$ 7
DiffEditor (Mou et al., 2024a) 2.34 $\pm$ 4
GoodDrag (Zhang et al., 2025) 5.86 $\pm$ 5
DragText (Choi et al., 2025) 2.73 $\pm$ 4
FastDrag (Zhao et al., 2024) 2.34 $\pm$ 5
@@
Ours 63.67 $\pm$ 16
Figure 5: Comparison between drag and move mode on Drag-Bench.
A total of 32 expert participants evaluated comparisons between methods on 32 cases randomly sampled from DragBench. For each comparison, method order positions were randomized and method identities were anonymized. Participants selected the preferred result according to predefined criteria (edit success, naturalness, and background preservation). Overall, LazyDrag was preferred in 61.88% of comparisons, outperforming all baselines (Tab. 4.4). More details are in Appendix A.4.
4.5 Comparison Between Drag and Move Modes
@@ 4.6 Ablation Study
We conduct an ablation study in which components are progressively removed from the full method. To keep functionality comparable when a component is absent, we adopt controlled replacements: (i) Without WTA and Latent Init (Sec. 3.2) we revert to latent warpage optimization of FastDrag (Zhao et al., 2024) as the latent initialization. (ii) Without ID Pres. and Attn Refine (Sec. 3.3) we switch to the attention-similarity matching and scaling introduced in CharaConsist (Wang et al., 2025b). Fig. 4.6 and Tab. 4.6 report benchmark results on Drag-Bench. Removing WTA and Latent Init increases MD and slightly reduces PQ and O, indicating that our initialization with the winner-takes-all fusion strategy and random initialization for inpainting regions suppresses repetitive artifacts and improves inpainting quality as proven in the figure. Further disabling background preservation causes additional drops in SC and O due to color shifting and artifacts in the background. Finally, replacing our correspondence-driven preservation with attention-similarity control leads to a sharp degradation, highlighting the sensitivity of full-strength inversion to mismatched attention alignment. The full method achieves the best performance.
@@
We conduct an ablation study on the effect of activation timesteps by varying the activation timestep to 20, 40, and 50, as shown in Fig. 7 and Tab. 4.6. From the results, we observe that increasing the number of the activation timestep leads to more accurate destination points for dragging, though it may introduce more warping artifacts. Conversely, reducing the activation timestep results in more natural outputs, but may cause slight variations in identity or motion. More results are given in Appendix B.5. For benchmark evaluations, we use 40 as a balanced value.
Figure 6: Qualitative cumulative ablation on Drag-Bench. Rows remove one component relative to the row above. When WTA and Latent Init are removed we use latent init in FastDrag. When ID Pres. and Attn Refine are removed we switch to CharaConsist attention-similarity control.
Figure 7: Qualitative ablation of activation timesteps on Drag-Bench. From left to right, the activation timestep is increased.
Table 3: Quantitative cumulative ablation on Drag-Bench under the same setting as Fig. 4.6
@@
Table 4: Quantitative ablation of activation timesteps on Drag-Bench.
@@
Additional results.
Appendix B presents additional evaluations on Drag-Bench, ablation studies with U-Nets, effects of text guidance, runtime analysis, and limitations.
5 Conclusion
@@
Ethics Statement
The development of advanced image editing technologies inevitably raises important ethical concerns. Although our method enhances editing precision through text and drag-based controls, it also introduces potential risks, including the creation of misleading or harmful visual content. To address this, we emphasize the importance of using such tools responsibly, with clear attention to transparency and user consent in practical deployments. In addition, the underlying pre-trained models may encode and reproduce societal biases, which could influence the outputs in unintended ways. We view this as an open research challenge and encourage future work aimed at bias detection and mitigation. All human evaluation participants were fully informed of the purpose of the study and provided consent before participation.
Reproducibility Statement
We have made every effort to ensure the reproducibility of LazyDrag. Detailed descriptions of the inference procedure and evaluation settings are provided in Sec. 3, Sec. 4.1 and Appendix A. All source code will be released to the public upon acceptance of this paper, enabling researchers to fully replicate and build upon our results.
@@ A.1 Inference Settings
EasyDrag (Hou et al., 2024) and CLIPDrag (Jiang et al., 2025) are excluded from comparison because their released implementations either fail to execute reliably or do not reproduce the results reported in the papers. DragGAN (Pan et al., 2023) is also excluded due to its inferior performance and slower processing speed compared to diffusion-based methods, as demonstrated in GoodDrag (Zhang et al., 2025).
@@
For our inversion process, we adopt the official inversion method of UniEdit-Flow (Jiao et al., 2025) but replace the editing component with our proposed strategy. We apply our correspondence-driven preservation (Sec. 3.3) only to the single-stream attention layers in FLUX.1 Krea-dev (Labs, 2025). Since additional manipulation in dual-stream attention layers does not lead to noticeable improvements (Deng et al., 2025; Yin et al., 2025b; Wang et al., 2025a), we adopt a more efficient and concise design by limiting modifications to single-stream layers only.
A.2 Implementation Details of Displacement Field Calculation
@@
$\lambda_{j}^{i}=\frac{\lVert\bm{v}_{j}^{i}\rVert_{2}}{\lVert\bm{d}_{i}\rVert_{2}}=\frac{\lVert\bm{p}_{j}-\bm{p}_{j}^{i}\rVert_{2}}{\lVert\bm{s}_{i}-\bm{e}_{i}\rVert_{2}}=\frac{\lVert\bm{p}_{j}-\bm{q}_{j}^{i}\rVert_{2}}{\lVert\bm{s}_{i}-\bm{q}_{j}^{i}\rVert_{2}}.$ (10)
@@
$\alpha_{j}^{i}=\begin{cases}\lVert\bm{p}_{j}-\bm{s}_{i}\rVert_{2}^{-1},&\bm{p}_{j}\neq\bm{s}_{i},\\ \infty,&\text{otherwise}.\end{cases}$ (11)
@@
$\bm{v}_{j}=\bm{v}_{j}^{i^{\star}}=\lambda_{j}^{i^{\star}}\,\bm{d}_{i^{\star}}.$ (12)
@@
$\bm{v}_{j}=\lambda_{j}^{i^{\star}}\,\bm{d}_{i^{\star}}+(\bm{r}-\mathbf{1})\otimes(\bm{p}_{j}-s_{i^{\star}}),$ (13)
@@
A.3 Evaluation Details
For the VIEScore evaluation, we follow GEdit-Bench (Liu et al., 2025b), using the same prompts for PQ and O. For SC, we adopt the instruction shown in Fig. 14, together with the source image, drag-instruction image, and the edited image. Score collection and calculation are carried out using the official GEdit-Bench codebase.
A.4 User Study Details
To evaluate the effectiveness of our method, we randomly selected 32 results for nine comparison methods on Drag-Bench (Shi et al., 2024b) and shuffled their indices to ensure a fair comparison. We invited 32 participants, each with relevant skills, to perform the tasks following the instructions provided through the user interface, as shown in Fig. 13.